# Detecting Harmful Memes with Decoupled Understanding and Guided CoT Reasoning

### v1

@@ 3.2.1 Guideline Crafting
Although harmfulness classification criteria across different contexts are disparate, we can find patterns of harmful content within each specific domain. These patterns can be summarized into fine-grained taxonomies of target or type. The target of a harmful meme is the entity at which the harmful message is directed, such as a vulnerable minority or protected group, an organization or a specific public figure, etc.The type of a harmful meme refers to a sub-category of harmfulness. For example, a misogynistic meme can be further divided into “stereotypes”, “body shaming”, “objectification”, “violence”, etc. Fersini et al. (2022). By prompting LLMs to first identify a potential target involved in a meme, or to classify whether a meme falls into a fine-grained type instead of the general “harmful” type, we can narrow the problem, reduce uncertainty, and guide the reasoning of LLMs to be more focused.
Examples & Exception
We articulate the patterns found in harmful memes in natural language as examples of harmful content, e.g., “Perpetuating stereotypes related to female’s domestic roles”, “Anti-semitic contents that make light of the Holocaust” etc.  We can further prompt LLMs to generate more examples and summarize them into more structured heuristics. These examples aim to encourage and support analogical reasoning of LLMs. Moreover, for a classification task, it is also important not to sacrifice precision solely for a high recall. Exception guidelines define the circumstances when a meme should be considered “harmless”, and may vary with datasets, usually reflecting the inherent bias and specificity in the original annotation.
3.2.2 CoT Prompting
[…]
4 Experiments
@@
We report the Accuracy (Acc) and macro-F1 (F1) score for performance evaluation in line with recent studies Mei et al. (2025); Huang et al. (2024). See Appendix C for implementation details.
Datasets
@@
We evaluate our proposed framework on seven widely used harmful meme detection datasets: (i) FHM(Kiela et al., 2020), (i) MAMI(Fersini et al., 2022), (i) PrideMM(Shah et al., 2024), (i) HarMeme(Pramanick et al., 2021a), (i) Harm-PPramanick et al., 2021b, (i) MultiOFF(Suryawanshi et al., 2020)and (i) GoatBench(Lin et al., 2024b). These datasets and benchmarks cover common harmful content in online memes, addressing different targets and topics across a broad range of sociocultural contexts. Table 5 summarizes the dataset statistics. More dataset details are in Appendix A.
@@
Pro-Cap RoBERTa 75.10 85.03 73.63
@@
ExplainHM GPT3.5+FLAN-T5 75.60 75.39 87.00 86.41 90.73 90.72
IntMeme FLAVA+RoBERTa 71.52 82.66 72.44
M3H-p-CoT Mistral-7B 80.28
@@
GPT-4o-mini 67.60 65.51 70.90 69.46 65.35 65.35 65.77 64.86 77.40 76.59 72.39 72.28
RGCL+Few-shot Qwen2VL-7B 63.50 68.10
@@
Ours: Unimodal Guided CoT
U-CoT+ Qwen2.5-14Bf 72.50 72.41 83.62 82.00 65.35 65.35 69.13 68.37 79.90 79.89 71.60 71.37
Mistral-12Bf 72.90 72.87 83.90 82.41 63.10 62.01 64.43 61.36 75.00 74.72 69.03 68.07
Table 1: Comparison with existing SOTA baselines under full-shot SFT or low-resource settings. The best SFT performance is in bold. The best zero-/few-shot performance is in red bold, with the second-best underlined. : LMM-RGCL Mei et al. (2025). RGCL+RKC/Few-shot is the out-of-distribution performance of the fine-tuned LMM-RGCL model with retrieval augmentation and few-shot examples, respectively.
Baselines
We consider the following baselines under two settings. For full-short supervised fine-tuning-based methods, we consider (i) LMM-RGCLMei et al. (2025), (ii) UMRYang et al. (2024), (iii) Pro-CapCao et al. (2023a), (iv) CapAlignJi et al. (2024), (v) MemeCLIPShah et al. (2024), (vi) M3H-p-CoTKumari et al. (2024), (vii) ExplainHMLin et al. (2024a), (viii) IntMemeHee and Lee (2025). For low-resource and training-free baselines, we include (ix) LoReHMHuang et al. (2024), (x) Mod-HATECao et al. (2024), (xi) GPT-4o-miniHurst et al. (2024). Since we extract LLMs’ prediction based on exact token match, the probability-based Area Under the Receiver Operating Characteristic Curve (AUC) score is not applicable. For baselines that evaluates fine-tuned classifiers on AUC, we only include their reported Acc results.
Model Scheme Hateful Harmful Misogyny Offensive
Acc F1 Acc F1 Acc F1 Acc F1
GPT-4V M-CoT 71.90 71.37 66.90 64.97 81.60 81.57 61.37 60.99
Qwen2.5 (14B) U-CoT 70.80 69.17 62.64 60.51 77.40 77.38 60.97 59.76
U-CoT+ 71.95 71.05 71.16 70.69 80.10 80.09 61.78 60.90
Mistral (12B) U-CoT 70.20 68.21 61.45 58.93 75.50 75.48 63.53 60.62
U-CoT+ 71.45 70.43 70.76 68.32 75.10 74.83 60.70 58.86
Table 2: Performance comparison of small LLMs against GPT-4V on the Goat-Bench dataset. GPT-4V results are given in Lin et al., 2024b. Our best performance is highlighted in bold.
LMMs & LLMs
We experiment with Llava1.6-7B Liu et al. (2023a) and Qwen2VL-7B Wang et al. (2024) as the representative small-scale LMMs for visual information extraction and generation of low-fidelity meme descriptions based on visual question answering. For small-scale LLMs, we evaluate the unimodal harmful meme detection performance of instruction-tuned Qwen2.5-14B Qwen et al. (2025), Mistral-12B Mistral (2024), Qwen2.5-7B Qwen et al. (2025), and Llama3.1-8B Grattafiori et al. (2024), as they represent state-of-the-art reasoning capabilities among open-source models with comparable parameter counts.
Table 1 presents the best zero-shot meme classification performance achieved by the best-performing Qwen2.5-14B and Mistral-12B models on the respective benchmarks, compared to a range of strong state-of-the-art baselines under both full-shot supervised fine-tuning and low-resource settings. The performance of Qwen2.5-7B and Llama3.1-8B are presented in Table 6. By default, each LLM reasons over visual information extracted exclusively by either Llava1.6-7B or Qwen2VL-7B. However, for the FHM dataset, we experiment with a specific collaborative setup between the two models: Qwen2VL-7B is responsible for extracting disability-related visual cues, while Llava1.6-7B handles the rest. This design is motivated by Qwen2VL-7B’s lower tendency to hallucinate when identifying such information. This FHM-specific configuration is denoted as “LLaVa+Qwen” in the subsequent illustrations.
We denote the traditional multimodal meme classification scheme based on LMMs under zero-shot CoT settings as M-CoT, where LMMs are directly prompted with the meme image along with the textual prompt. Our proposed unimodal guided CoT-based harmful meme detection scheme is denoted as U-CoT+, while U-CoT denotes the ablation variant, in which the LLMs reason solely based on meme descriptions without referring to our human-crafted guidelines.
@@
Qwen2.5-14Bf Qwen2VL-7B U-CoT $\uparrow$70.10 $\uparrow$70.02 61.02 57.92 $\uparrow$60.00 $\uparrow$59.91 53.69 53.15 $\uparrow$75.10 $\uparrow$75.10 68.44 68.42
LLaVa1.6-7B $\uparrow$68.00 $\uparrow$67.70 59.60 51.46 $\uparrow$63.94 $\uparrow$63.84 $\uparrow$64.43 $\uparrow$62.93 $\uparrow$76.40 $\uparrow$76.38 $\uparrow$66.47 $\uparrow$66.47
LLaVa+Qwen $\uparrow$69.10 $\uparrow$68.83
Qwen2VL-7B U-CoT+ $\uparrow$72.50 $\uparrow$72.41 $\uparrow$81.92 $\uparrow$81.00 $\uparrow$65.35 $\uparrow$65.35 $\uparrow$63.09 $\uparrow$62.41 $\uparrow$78.60 $\uparrow$78.59 $\uparrow$70.41 $\uparrow$70.04
LLaVa1.6-7B $\uparrow$71.50 $\uparrow$71.48 $\uparrow$83.62 $\uparrow$82.00 63.94 63.67 $\uparrow$69.13 $\uparrow$68.37 $\uparrow$79.90 $\uparrow$79.89 $\uparrow$71.60 $\uparrow$71.37
LLaVa+Qwen $\uparrow$72.00 $\uparrow$71.98
Mistral-12Bf Qwen2VL-7B U-CoT $\uparrow$68.30 $\uparrow$68.09 65.54 $\uparrow$63.48 $\uparrow$58.87 $\uparrow$58.32 65.10 62.53 $\uparrow$73.00 $\uparrow$72.99 62.72 62.16
LLaVa1.6-7B $\uparrow$65.30 $\uparrow$64.60 62.43 54.60 $\uparrow$58.03 $\uparrow$56.53 $\uparrow$61.74 56.89 $\uparrow$76.40 $\uparrow$76.38 $\uparrow$62.33 $\uparrow$61.68
LLaVa+Qwen $\uparrow$67.20 $\uparrow$66.70
Qwen2VL-7B U-CoT+ $\uparrow$71.20 $\uparrow$71.19 $\uparrow$83.90 $\uparrow$82.41 $\uparrow$63.10 $\uparrow$62.01 59.06 57.33 $\uparrow$74.90 $\uparrow$74.81 $\uparrow$69.03 $\uparrow$68.07
LLaVa1.6-7B $\uparrow$70.00 $\uparrow$69.88 $\uparrow$81.07 $\uparrow$78.57 $\uparrow$59.72 56.40 $\uparrow$64.43 $\uparrow$61.36 75.00 74.72 $\uparrow$68.05 $\uparrow$67.82
LLaVa+Qwen $\uparrow$72.90 $\uparrow$72.87
Table 3: Ablation studies comparing different prompting schemes for zero-shot harmful meme detection. Performance improvements from multimodal (M-CoT) to unimodal text-only (U-CoT) settings are highlighted by $\uparrow$, while improvements from plain to guided CoT (U-CoT+) are highlighted by $\uparrow$. The best unimodal results are in bold.
4.1 U-CoT+ vs. Supervised Fine-tuned Baselines
As shown in Table 1, although performance gaps remain between zero-shot guided CoT prompted small-scale LLMs and state-of-the-art fully supervised fine-tuned models, however, on FHM, HarMeme, MAMI and PrideMM, their zero-shot classification performance is nearly comparable to that of SOTA full-shot SFT baselines. For example, Mistral-12B achieves 72.90 (Acc) and 72.87 (F1) on FHM, which has surpassed IntMeme and come within 3 points of the performance achieved by ProCap and ExplainHM. Notably, its performance on HarMeme (83.90/82.41) is comparable to most SFT baselines, surpassing IntMeme and CapAlign, achieving a notably higher macro-F1. Both LLMs perform well on the MAMI dataset, with Qwen2.5-14B achieving an accuracy of 79.90 and a macro-F1 score of 78.90, which is on par with the performance of LMM-RGCL. While slightly lower, the best zero-shot performance of small LLMs on the challenging PrideMM dataset remains comparable to that of SFT baselines, with a performance gap within 7 points.
The performance disparity between small LLMs and fully supervised baselines is most pronounced on datasets of U.S. political memes. Fully supervised models achieve similarly high accuracy on Harm-P, surpassing 90%, while small LLMs struggle to exceed 66%. Although MultiOFF and Harm-P share similar contexts and meme styles, SFT baselines do not exhibit equally strong performance on MultiOFF. Specifically, UMR and LMM-RGCL achieve F1 scores of 69.96 and 64.80, respectively, making the zero-shot performance of Qwen2.5-14B (69.13/68.37) competitive by comparison. The performance drop of SFT baselines on MultiOFF is likely due to the reduced amount of training data, as shown in Table 5, with only 445 instances which are significantly fewer than the 2,939 available in Harm-P. This highlights the inherent limitations of supervised fine-tuning methods under low-resource conditions, where insufficient training data leads to degraded performance.
Model Scheme FHM HarMeme MAMI
Acc F1 Acc F1 Acc F1
[…]
Qwen2.5 (14B)
U-CoT 70.10 70.02 61.02 57.92 76.40 76.38
U-CoT+L $\uparrow$70.50 $\uparrow$70.49 $\uparrow$67.23 $\uparrow$65.90 $\uparrow$76.70 $\uparrow$76.70
U-CoT+ 72.50 72.41 83.62 82.00 79.90 79.89
Mistral (12B)
U-CoT 68.30 68.09 65.54 63.48 76.40 76.38
U-CoT+L 65.80 63.78 65.54 59.15 67.50 65.24
U-CoT+ 72.90 72.87 83.90 82.41 75.00 74.72
[…]
4.2 U-CoT+ vs. Low-Resource Baselines
Although operating under a zero-shot setting, low-resource baselines with retrieval-based augmentation such as LoReHM and LMM-RGCL+RKC, further benefit from a majority voting mechanism that leverages the ground-truth labels of the top-K most similar memes retrieved in the embedding space. However, the small-scale LLMs in our proposed framework rely solely on human-crafted guidelines, without the need for any auxiliary ground-truth signals or few-shot examples for additional reference. Notably, the performance of the supervised fine-tuned LMM-RGCL degrades significantly in out-of-distribution zero-shot inference settings, even when supported by the retrieval-based KNN classifier (RKC).

### v2

@@ 3.2.1. Guideline Crafting
  • Fine-grained Taxonomy Context-specific harmful memes often have fine-grained taxonomies of target or type. A harmful meme’s target is the entity toward which its harmful message is directed, such as a vulnerable minority, protected group, organization or specific public figure, etc.  The type of a harmful meme refers to the sub-category of harmfulness it falls in, such as “stereotypes” or “body shamingetc., in misogynistic memes (Fersini et al., 2022). Based on such taxonomies, by first prompting LLMs to identify a meme’s potential target or breaking down into fine-grained classification sub-tasks, this principle seeks to narrow the problem and reduce uncertainty, guiding more focused reasoning through the use of target- or type-specific guidelines.
  • Patterns This principle aims to foster analogical reasoning by prompting LLMs with distilled patterns of harmful memes e.g., “Perpetuating stereotypes on female’s domestic roles”, “Anti-Semitic contents that make light of the Holocaust” etc.To derive such patterns with further reduced human efforts, we can leverage LLMs’ in-context learning abilities to extend and integrate human-provided examples into structured heuristics.
  • Exception When detecting harmful memes, it is crucial not to prioritize high recall at the cost of precision. This principle aims to prevent over-censorship by introducing exception cases in which memes should be considered “harmless” despite containing traces of potentially sensitive content. These exceptions may vary across contexts and domains, often reflecting specific evaluation criteria or inherent biases in data annotation.
3.2.2. Zero-shot CoT Prompting
Following the widely adopted implementation of chain-of-thought (CoT) reasoning (Wei et al., 2022), we prompt unimodal LLMs in a zero-shot setting using a heuristic prompt $P^{CoT}$ such as “Now, let’s think step by step” or “Now, let’s apply the guidelines one by one”. This encourages LLMs to produce the final classification decision $CLS$ through detailed step-by-step rationales $R$ by reasoning over the high-fidelity meme description $D_{h}$ together with our human-crafted harmful meme detection guidelines $GL$. Consequently, we have: $CLS,R=\mathbf{LLM_{\text{U-CoT+}}}(D_{h},GL,P^{CoT})$.
4. Experiments
@@
We report the Accuracy (Acc.) and macro-F1 (F1) score for performance evaluation in line with recent studies (Mei et al., 2025; Huang et al., 2024). As our method extracts LLM predictions based on exact token match, probability-based AUC scores are not applicable. We report the corresponding accuracy results for baselines that evaluate using AUC. See Appendix B for implementation details.
@@
  1. Datasets. We evaluate our proposed framework on seven widely used harmful meme detection datasets: (i) FHM(Kiela et al., 2020), (ii) HarMeme(Pramanick et al., 2021a), (iii) Harm-P(Pramanick et al., 2021b), (iv) MultiOFF(Suryawanshi et al., 2020), (v) MAMI(Fersini et al., 2022), (vi) PrideMM(Shah et al., 2024)and (vii) GoatBench(Lin et al., 2024b). These datasets and benchmarks cover various types of harmfulness against diverse targets across a wide range of sociocultural contexts.
  2. Baselines. For full-shot SFT methods, we consider: (i) LMM-RGCL(Mei et al., 2025), (ii) UMR(Yang et al., 2024), (iii) ExplainHM(Lin et al., 2024a), (iv) Pro-Cap(Cao et al., 2023a), (v) IntMeme(Hee and Lee, 2025), (vi) CapAlign(Ji et al., 2024), (vii) M3Hop-CoT(Kumari et al., 2024), (viii) MemeCLIP(Shah et al., 2024). For low-resource training-free methods, we include (i) Zero-shot inference with GPT-4o-mini(OpenAI, 2024), (ii) LoReHM(Huang et al., 2024), (iii) MInD(Liu et al., 2025), (iv) Mod-HATE(Cao et al., 2024), and (v) GPT-4V(OpenAI, 2023).
  3. Models. We experiment with representative state-of-the-art lightweight open-source models, including LLaVA1.6-7B (Liu et al., 2023a) and Qwen2VL-7B(Qwen2-VL, 2024) as LMMs, and instruction-tuned Qwen2.5-14B and 7B (Report, 2025), Mistral-12B (Mistral, 2024), and Llama3.1-8B (AI@Meta, 2024) as unimodal LLMs.
Table 1. Comparison with SOTA SFT and low-resource baselines. The best SFT results are in bold. The best zero-/few-shot performance is highlighted in bold and underlined; second-best is underlined. : RGCL+RKC/FS is the out-of-distribution performance of LMM-RGCL (Mei et al., 2025) with retrieval augmentation and few-shot examples, respectively.
@@
ExplainHM GPT3.5+FLAN-T5 75.60 75.39 87.00 86.41 90.73 90.72
Pro-Cap BLIP-2+RoBERTa 75.10 85.03 73.63
IntMeme FLAVA+RoBERTa 71.52 82.66 72.44
@@
Zero-shot GPT-4o-mini 67.60 65.51 70.90 69.46 65.35 65.35 65.77 64.86 77.40 76.59 72.39 72.28
RGCL+FS Qwen2VL-7B 63.50 68.10
@@
Ours: Unimodal Guided Zero-shot CoT
[…]
Table 2. Performance on Goat-Bench. S.: Scheme. M: M-CoT; U: U-CoT; U+: U-CoT+. GPT-4V results are reported in (Lin et al., 2024b). Our best results are highlighted in bold.
Model S. Hateful Harmful Misogyny Offensive
Acc. F1 Acc. F1 Acc. F1 Acc. F1
GPT-4V M 71.90 71.37 66.90 64.97 81.60 81.57 61.37 60.99
Qwen2.5-14B U 70.80 69.17 62.64 60.51 77.40 77.38 60.97 59.76
U+ 71.95 71.05 71.16 70.69 80.10 80.09 61.78 60.90
Mistral-12B U 70.20 68.21 61.45 58.93 75.50 75.48 63.53 60.62
U+ 71.45 70.43 70.76 68.32 75.10 74.83 60.70 58.86
Table 1 and Table 2 compare our best zero-shot harmful meme detection performance achieved by the best-performing unimodal LLMs under U-CoT+ to all the state-of-the-art (SOTA) baselines under either full-shot SFT or low-resource settings. We denote the traditional LMM-based multimodal classification scheme under zero-shot CoT as M-CoT, where the LMM reasons over the prompt and image. U-CoT denotes U-CoT+’s ablation variant, where the LLM performs zero-shot CoT reasoning over textual inputs but without referring any supportive information such as guidelines.
Table 3. Ablation studies. The best low-resource results (with open-source models, without ensemble) are in bold, with the second-best underlined. $\Delta$: Performance gain over 0-shot guideline-free U-CoT. $\bar{\Delta}_{\textit{U-CoT+}}$: The average performance gain of U-CoT+ under guideline perturbations. $\Delta_{\textit{U-CoT+L}}$: The performance gain of U-CoT+ with LoReHM insights. $\bar{\Delta}_{\textit{U-CoT+FS}}$: The average performance gain of U-CoT+ with few-shot (4,6,8 and 10-shot) examples.
@@
Qwen2.5-14B 7B LMM U-CoT 70.10 70.02 59.60 51.46 60.00 59.91 64.43 62.93 76.40 76.38 66.47 66.47
U-CoT+ 72.50 72.41 83.62 82.00 65.35 65.35 69.13 68.37 79.90 79.89 71.60 71.37
$\bar{\Delta}_{\textit{U-CoT+}}$ 1.78 1.75 22.11 28.58 4.22 4.23 0.50 0.98 1.96 1.93 3.70 3.60
$\Delta_{\textit{U-CoT+L}}$ 0.40 0.47 7.63 14.44 0.30 0.32
$\bar{\Delta}_{\textit{U-CoT+FS}}$ -0.70 -0.62 4.81 1.78 1.20 1.21 -0.67 -1.77 1.25 1.25 0.44 0.26
7B LMMs Ensemble 72.40 72.23 82.77 82.13 65.07 65.05 62.42 62.21 79.90 79.85 $\uparrow$73.37 $\uparrow$73.36
GPT-4o-mini U-CoT+ 73.00 73.00 83.90 83.05 62.25 62.25 60.40 60.22 79.10 78.78 68.44 68.39
Mistral-12B 7B LMM U-CoT 67.20 66.70 65.54 63.48 58.87 58.32 61.74 56.89 73.00 72.99 62.72 62.16
U-CoT+ 72.90 72.87 83.90 82.41 63.10 62.01 64.43 61.36 74.90 74.81 69.03 68.07
$\bar{\Delta}_{\textit{U-CoT+}}$ 4.22 4.70 14.76 14.94 4.51 4.04 2.69 4.43 2.20 2.09 6.46 6.49
$\Delta_{\textit{U-CoT+L}}$ -1.40 -2.92 0.00 -4.33 -5.50 -7.75
$\bar{\Delta}_{\textit{U-CoT+FS}}$ -0.73 -0.96 1.20 -4.78 1.77 -1.74 -0.67 -5.70 1.15 0.77 0.79 -1.41
7B LMMs Ensemble $\uparrow$73.50 $\uparrow$73.43 83.90 82.86 63.10 62.16 63.09 61.53 $\uparrow$76.00 $\uparrow$75.61 $\uparrow$70.02 $\uparrow$70.02
GPT-4o-mini U-CoT+ 71.10 71.09 81.92 80.61 62.82 61.27 67.79 66.12 75.20 74.84 66.27 66.21
Llama3.1-8B 7B LMM U-CoT 63.50 61.44 62.43 50.64 62.25 58.78 63.76 57.55 68.70 68.10 60.55 58.49
U-CoT+ 68.00 66.98 71.19 61.44 61.13 56.21 68.46 62.35 73.50 73.50 64.50 63.26
7B LMMs Ensemble $\uparrow$69.50 $\uparrow$69.00 $\uparrow$75.14 $\uparrow$69.34 60.56 56.29 67.79 61.79 $\uparrow$74.00 $\uparrow$74.00 $\uparrow$67.85 $\uparrow$67.58
Qwen2.5-7B 7B LMM U-CoT 66.70 66.45 59.89 59.82 58.59 57.46 63.76 63.72 70.60 70.59 63.51 62.38
U-CoT+ 67.50 67.50 77.12 75.26 62.54 60.20 67.11 66.63 74.40 74.40 67.26 66.90
7B LMMs Ensemble 67.00 66.98 70.34 70.23 $\uparrow$62.82 $\uparrow$61.76 63.09 63.09 $\uparrow$75.80 $\uparrow$75.78 67.06 67.03
  • Note: U-CoT results outperforming M-CoT are blue. U-CoT+ results based on $\leq$8B LLMs that outperform M-CoT are green. $\uparrow$: Ensemble results further improving on U-CoT+.
4.1. Comparing U-CoT+ to Baselines
U-CoT+ vs. Supervised Fine-tuned Baselines As shown in Table 1, performance gaps remain between the state-of-the-art fully supervised fine-tuned baselines and lightweight LLMs under our proposed unimodal guided zero-shot CoT prompting, but they have narrowed to a generally acceptable and promising range, especially given our focus on low-resource settings. Specifically, LLMs’ zero-shot performance has become nearly comparable to that of full-shot baselines on FHM, HarMeme, MAMI and PrideMM. On FHM, Mistral-12B outperforms IntMeme and comes within 3 points of the performance of ProCap and ExplainHM. Notably, its performance on HarMeme is highly comparable to that of most SFT baselines, surpassing IntMeme and CapAlign, even with better macro-F1. Moreover, both LLMs perform well on MAMI, with Qwen2.5-14B matching the performance of LMM-RGCL. Its best zero-shot result on the challenging PrideMM also falls within 7 points to that of the two SFT baselines. On the U.S. political meme dataset Harm-P, SFT methods consistently achieve high accuracy over 90%, whereas zero-shot inference, even with GPT-4o-mini, struggles to exceed 70%, which reflects the complex, highly subjective, heterogeneous, and target-dependent nature of defining harmfulness in Harm-P. In contrast, SFT methods perform less strongly on MultiOFF, despite its contextual and stylistic similarity to Harm-P, making Qwen2.5-14B’s zero-shot performance comparable. SFT’s decline on MultiOFF suggests a drawback of reduced training data (455 vs. 2.9K in Harm-P). This observation verifies SFT methods’ heavy reliance on large-scale training data, revealing their inherent limitations under low-resource conditions.